OpenStudy (anonymous):

(1-sin^2x)(1+sin^2x) = 2cos^2x - cos^4x

OpenStudy (john_es):

Use \[\cos^2x=1-\sin^2x\]Then, \[cos^2x(2+\cos^2x)\]Multiply it and you'll find the way.

OpenStudy (anonymous):

for which side?

OpenStudy (john_es):

I begin from the left to reach the right side.

OpenStudy (anonymous):

so your saying, (cos^2x)( 1+sin^2x) on the left side?

OpenStudy (john_es):

Exact.

OpenStudy (anonymous):

Then what?

OpenStudy (john_es):

I'll write it again because I miss a sign :)

OpenStudy (john_es):

\[(1-\sin^2x)(1+\sin^2x) =\cos^2x(1+1-\cos^2x)=\cos^2x(2-\cos^2x)=\\ =2\cos^2x-\cos^4x\]

OpenStudy (john_es):

That's all ;)

OpenStudy (anonymous):

They don't equal though?

OpenStudy (john_es):

Well, the line I wrote you, demonstrate the equality you wrote in the problem. Do you see it? It begins from the left side you wrote and finish in the right side.

OpenStudy (anonymous):

ok how did you get (1+sin^2x) to (1+1 - cos^2x)

OpenStudy (john_es):

Well, we can apply this identity, \[\sin^2x+\cos^2x=1\]Then \[\sin^2x=1-\cos^2x\]And substitute.

OpenStudy (anonymous):

Oh I see it now, I didn't now you can do that with the + 1 with it

OpenStudy (john_es):

Apply it when you need it ;), it is a very tricky identity.

OpenStudy (anonymous):

I got another one

OpenStudy (anonymous):

secx/tanx - tanx/secx = cosx cotx

OpenStudy (john_es):

Well, we need to use, \[\sec x=1/\cos x\]\[\tan x=\sin x/\cos x\]\[\cot x=\cos x/\sin x\]Then \[(1/\cos x)/(\sin x/\cos x)-(\sin x/\cos x)/(1/\cos x)=\frac{1}{\sin x}-\sin x=\\ =\frac{1-\sin^2x}{\sin x}=\frac{\cos^2x}{\sin x}=\cos x\frac{\cos x}{\sin x}=\cos x\cot x\]

OpenStudy (anonymous):

alright I think found another way. I times secx/tanx by secx/secx and tanx/secx by tanx/secx, which leads to sec^2x - tan^2x on top which is 1. So 1 over tanx equeal cotx. and 1 over secx equals cosx

OpenStudy (john_es):

I would say you cannot do that, because you are not multiplying by 1 each fraction. So you, finally, would not obtain the equality you are searching for.

OpenStudy (anonymous):

It worked out perfectly for me, so Im keeping it

OpenStudy (john_es):

Okay, so when you finally sum the fractions do you obtain cos x* cot x?

OpenStudy (john_es):

If this the case then something wrong is happening in the Universe :)

OpenStudy (anonymous):

yeah 1 over secx = cosx and 1over tanx = cotx

OpenStudy (anonymous):

I got one more identity, -1/tanx - secx = 1+sinx/cosx

OpenStudy (john_es):

Are you serious?

OpenStudy (anonymous):

yes?

OpenStudy (john_es):

\[-\frac{1}{\tan x}-\sec x=-\frac{1}{\sin x/\cos x}-1/\cos x=-\frac{\cos x}{\sin x}-\frac{1}{\cos x}=\\ =-\frac{\sin x+\cos^2x}{\sin x\cos x}\]

OpenStudy (john_es):

Your last relation is impossible. You only need to do x=0 and the relation doesn't hold.

OpenStudy (anonymous):

Oh alright, I can do this one on my own. You helped me enough! Helping a stranger to do math, is a great deed. Thank you very much :)

OpenStudy (john_es):

Don't worry. But in case you need, I think wolfram alpha could be very helpful to check this relations. I check the last one with it, using x=1. http://www.wolframalpha.com/input/?i=-1%2FTan%5B1%5D-1%2FCos%5B1%5D%3D%3D1%2BSin%5B1%5D%2FCos%5B1%5D

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You have the identity wrong

OpenStudy (john_es):

Could you please explain me why?

OpenStudy (anonymous):

Let me re write it?

OpenStudy (john_es):

Please

OpenStudy (anonymous):

-1/tanx - secx = 1 +sin x/cos x

OpenStudy (john_es):

And what is the difference between this and mine?

OpenStudy (john_es):

\[\sec x=\frac{1}{\cos x}\]

OpenStudy (anonymous):

Instead of sec x you had cos x

OpenStudy (john_es):

Well, I think you note that, \[\sec x=\frac{1}{\cos x}\]

OpenStudy (john_es):

But don't worry, Wolfram Alpha also know about Secx, so here is the link with the same result, http://www.wolframalpha.com/input/?i=-1%2FTan%5B1%5D-Sec%5B1%5D%3D%3D1%2BSin%5B1%5D%2FCos%5B1%5D

OpenStudy (john_es):

Do you see it? The test is false, so the relation doesn't hold.

OpenStudy (anonymous):

Maybe my teacher made this one wrong on purpose

OpenStudy (anonymous):

I see the relation

OpenStudy (john_es):

May be your teacher put it wrong, it is strange, but it doesn't hold. Sorry.

OpenStudy (anonymous):

Alright well your the expert, so it has to be wrong

OpenStudy (john_es):

May be a sign or something. I have tried all the combinations but I didn't find anything that your teacher cpuld put as a problem with this expression.

OpenStudy (anonymous):

I got the input to say true when I wrote it

OpenStudy (anonymous):

-1/(tanx - secx) = (1+sinx) / cosx

OpenStudy (anonymous):

I forgot the parentheses